OpenStudy (anonymous):

Find the point on the parabola 4y= x^2 nearest to the point (1,2).

OpenStudy (amistre64):

we need to find a normal to the line that perps to 1,2 then

OpenStudy (amistre64):

first; lets define the slope of the line for any given point on it: 4y = x^2 4 y' = 2x y' = (1/2)x this is tell us the slope of the tangent line at any given x

OpenStudy (anonymous):

Can i answer through skype?? If not answer is (3.538584704/2,3.538584704^2/8)

OpenStudy (amistre64):

to perp it; we flip and negate it

OpenStudy (amistre64):

slope = (1/2)x perp slope = -2/x and we want it to run thru the point (1,2); and hopefully im thinking thru this correctly :)

OpenStudy (amistre64):

the equation for the the perp line becomes: y = -2x/x + 2(1)/x + 2 y = -2 + 2/x + 2 y = 2/x .... i got me doubts

OpenStudy (amistre64):

so where does: 4y= x^2 and y = 2/x meet by chance? x^2/4 = 2/x x^3 = 8 x=2 hmmmm if im right, and i doubt it; id say at (2,1)

OpenStudy (zarkon):

I get (2,1) also

OpenStudy (zarkon):

you can also use the distance formula to solve this problem

OpenStudy (amistre64):

well, if it isnt right, then at least its a consensus lol

OpenStudy (zarkon):

it is correct

OpenStudy (nikvist):

\[d^2=(x-1)^2+(y-2)^2=(x-1)^2+(x^2/4-2)^2\]\[(d^2)'=2(x-1)+2(x^2/4-2)\frac{x}{2}=0\]\[2x-2+\frac{x^3}{4}-2x=0\quad\Rightarrow x^3=8\quad,\quad x=2\quad,\quad y=1\]